OpenStudy (anonymous):

Two functions, A and B, are described as follows: Function A y = 8x + 3 Function B The slope is 1 and the y-intercept is 4. How much more is the slope of function A than the slope of function B? 1 7 8 9

OpenStudy (anonymous):

@iGreen @AnswerMyQuestions

OpenStudy (anonymous):

plz help

OpenStudy (anonymous):

The equation is y = mx + b, where m is the slope, so can you find the slope for equatioon 1?

OpenStudy (igreen):

y = 8x + 3 This is in the form of y = mx + b, where 'm' is the slope. So what's the slope here?

OpenStudy (igreen):

*equation @AnswerMyQuestions

OpenStudy (anonymous):

:P

OpenStudy (anonymous):

8 is the slope

OpenStudy (igreen):

Yes!

OpenStudy (igreen):

And the slope of the 2nd one is 1. So subtract 8 - 1.

OpenStudy (anonymous):

7

OpenStudy (igreen):

Yes, so the slope for function A is 7 more than the slope for function B.

OpenStudy (anonymous):

the slope for b is 8 so the awnser is A(1)

OpenStudy (anonymous):

No.

OpenStudy (anonymous):

You subtract 8 - 1 to get your answer.

OpenStudy (anonymous):

oh so it is B (7)

OpenStudy (anonymous):

is that right

OpenStudy (igreen):

Yes, it's 7.

OpenStudy (anonymous):

thank you